OpenStudy (anonymous):

Write the expression 2 to the 5/3 power in radical form.

OpenStudy (anonymous):

\[(\sqrt[3]{2})^{5}\]

OpenStudy (anonymous):

Bottom number outside the sqrt sign and top number is the exponent of the problem

OpenStudy (anonymous):

Do you understand?

OpenStudy (anonymous):

okay thanks!!

OpenStudy (anonymous):

np

OpenStudy (anonymous):

if the radical is \[\sqrt[6]{10}\] then is the exponent \[10^{1/6}\] ?

OpenStudy (anonymous):

Correct

OpenStudy (anonymous):

if it is \[(\sqrt[2]{3})^{5}\] it would be \[3^{\frac{ 5 }{ 2 }}\]

OpenStudy (anonymous):

alright thanks again

OpenStudy (anonymous):

np